OpenStudy (anonymous):

y = 2x + 4 2x - y = 6 Solve the system of equations using substitution.

OpenStudy (anonymous):

from the first part y= -4-x x-(-4-x)=2 x+4+x=2 2x=-2 x=-1 plug in y= -4-(-1) y=-4+1 y= -3 hence x= -1 and y= -3